OpenStudy (kenialove):

A cone has a height of 24 inches , a slant height of 25 inches , and a diameter of 14 inches. What is the surface area of the cone ? A. 1,176 pi^2 B. 392 pi^2 C. 224 pi^2 D. 178 pi^2 I know how to plug in the numbers and everything but I forgot the formula to find my answer .. Can someone help me ?

jimthompson5910 (jim_thompson5910):

The surface area of a cone is \[\Large SA = \pi*r^2 + \pi*r*s\] SA = surface area r = radius s = slant height

jimthompson5910 (jim_thompson5910):

In this problem, r = 7 since the diameter is 14 and you cut the diameter in half to get the radius. Also, s = 25 is the slant height.

OpenStudy (kenialove):

Thanks so much

OpenStudy (kenialove):

So it would be .. A = 3.14 x 7^2 + 3.14 x 7 x 25

OpenStudy (kenialove):

I got the answer 703.36 but that's not one of the choices in the question .. Can u help me ? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

leave pi as it is

jimthompson5910 (jim_thompson5910):

think of pi as a variable if you have to

OpenStudy (kenialove):

Okay I'll try that

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (kenialove):

I think I need help cause I'm getting a bigger number

jimthompson5910 (jim_thompson5910):

what is 7^2 equal to?

OpenStudy (kenialove):

49

jimthompson5910 (jim_thompson5910):

how about 7*25, what is that equal to?

OpenStudy (kenialove):

175

jimthompson5910 (jim_thompson5910):

So... \[\Large SA = \pi*r^2 + \pi*r*s\] \[\Large SA = \pi*7^2 + \pi*7*25\] \[\Large SA = \pi*49 + \pi*175\] \[\Large SA = 49\pi + 175\pi\] \[\Large SA = 224\pi\]

jimthompson5910 (jim_thompson5910):

notice how I added 49pi to 175pi to get 224pi similar to me saying 49x+175x = 224x

OpenStudy (kenialove):

Ohhh I know where I missed up Thanks so much

jimthompson5910 (jim_thompson5910):

glad to be of help